Calculate the crop growth rate if the dry weight of the corn crop was 14g on 18 July, 2019 and dry weight of corn crop was 35g obtained on 02 August 2019.
1
Expert's answer
2020-08-10T19:53:35-0400

By definition, the growth rate is:


"G = \dfrac{m_f - m_i}{\Delta t}"

where "m_f = 35g" is the final dry weight of corn crop, "m_i = 14g" is the initial dry weight of corn crop, "\Delta t" is the number of days passed since the starting moment of time (since 18 July, 2019). Using the calendar we can conclude, that "\Delta t = 20 \space days". Hence, obtain:


"G = \dfrac{35g - 14g}{20\space days} = 1.05\space \dfrac{g}{day}"

Answer. 1.05 g/day.